# Repeating Flares, X-ray Outbursts and Delayed Infrared Emission: A Comprehensive Compilation of Optical Tidal Disruption Events

### v1

@@ B.3.3 Spectral Extraction
The source spectra were extracted with the evselect task from circular regions centred at the X-ray source coordinates with radii equal , while background spectra were extracted from annuli centred at the X-ray source coordinates, with inner and outer radii equal to and , respectively, excluding nearby detected sources. The corresponding arf and rmf files were generated with the rmfgen and arfgen tasks to take into account time and position-dependent EPIC responses. Again, inner regions of high pileup were estimated and excluded using the epatplot task through the distortion of pattern distribution, following the procedure explained in the SAS Data Analysis Threads212121https://www.cosmos.esa.int/web/xmm-newton/sas-thread-epatplot.
@@ B.4 X-Ray Spectral Fitting
Spectral fitting was performed with the Sherpa222222http://cxc.harvard.edu/sherpa modelling and fitting application (Freeman et al. 2001) in the energy range for Swift-XRT and XMM-Newton-EPIC spectra, and in the energy range for Chandra-ACIS spectra, adopting Gehrels weighting (Gehrels 1986). Source spectra were binned to a minimum of 20 counts/bin to ensure the validity of statistics. In addition, for the EPIC spectra we excluded from the spectral fitting the band due to variable Al K lines, and fitted simultaneously MOS1, MOS2 and PN spectra, if available.
For the spectral fitting we used two different models:
1.
a model comprising an absorption component fixed to the Galactic value and a power-law with slope
2.
a model comprising an absorption component fixed to the Galactic value and a black body with temperature
In order to select the best fit model we made use of the Akaike information criterion (AIC, see for example Liddle 2007):
where is the number of model parameters, so we selected as best fit model the one that provides the lower value of AIC.
B.5 X-ray Catalogs
@@ B.5.1 XMM Catalogue
We cross-matched the 4XMM-DR13 catalog with our catalogue of TDEs using a search radius. Since the fluxes reported in 4XMM-DR13 are derived from the count-rates assuming a model comprising a power-law with slope 1.7 and a Galactic absorption of (Webb et al. 2020), for uniformity we converted the count-rates for each detector to intrinsic fluxes using PIMMS232323https://heasarc.gsfc.nasa.gov/docs/software/tools/pimms.html, assuming a model comprising an absorption component fixed to the Galactic value and a black body with a temperature of . The fluxes for each detector and their respective uncertainties were there converted in mean flux and uncertainty using the same procedure illustrated in Sect. B.3.2.
@@ B.5.2 EROSITA
We cross-matched the eRASS1 catalogue with our catalogue of TDEs using a search radius The fluxes reported in eRASS1 are derived from the count-rates assuming a model comprising a power-law with slope 2.0 and a Galactic absorption of (Merloni et al. 2024), therefore we converted the count-rates reported in the catalogue to intrinsic fluxes using PIMMS, again assuming a model comprising an absorption component fixed to the Galactic value and a black body with a temperature of .
For the TDE sources for which we did not find an eRASS1 counterpart, we obtained eRASS1 upper limits242424https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/UpperLimitServer_dr1/. The upper limit eRASS fluxes were again converted in fluxes using PIMMS.
@@ Appendix C Archival repeating flares
C.1 AT 2022dbl
This source was proposed to be a pTDE by Lin et al. (2024). The first flare was observed in 2022, while the second in 2024, with a 1.9 year interval. The light curve for this TDE is plotted in the upper panel of Fig. 6(c). Lin et al. (2024) argue that due to the similar optical spectra of the two flares, more specifically the broad Balmer, N iii and possible He ii emission lines, the most probably scenario is the repeating pTDE one. We also compare the shape of the two flares in the bottom panel of Fig. 6(c) and find that they are similar.
Moreover, Lin et al. (2024) tried to structure the orbit of the star around the SMBH and found, under the assumption of a BH of mass equal to , an elliptical orbit with period 710 days (time interval between the two flares) and a solar-like star, the eccentricity to be 0.997. Nonetheless, the pTDE nature of this phenomenon will be confirmed if a third flare is observed, since the orbital period of this event is calculated to be relatively short (1.9 years).
C.2 AT 2022agi
As briefly mentioned in Sect. 2, this TDE was studied by Sun et al. (2024), showcasing two flares. The first flare was observed in 2010, while the second in 2021, with a time interval of years (Sun et al. 2024). The light curve for this source is plotted in the upper panel of Fig. 6(d). Similarly to AT 2024pvu, the first flare was observed by the CRTS survey. The second outburst was observed by both the ATLAS and Gaia surveys, in multiple wavebands. Sun et al. (2024) discuss in great detail the nature of the repeating flares; whether it is a pTDE, double TDE (binary star system disrupted by the SMBH) or two separate events. In the end they are not able to exclude any of the aforementioned categories, with all being theoretically possible.
C.3 AT 2021mhg
This source was first discovered to be a TDE spectroscopically in 2021, while it was later reported to have a re-brightening in 2023 (Munoz-Arancibia et al. 2023). However, Somalwar et al. (2023b) attribute the second flare to a supernova (SN) Type Ia, that happened at the same position as the original 2021 TDE (see their Appendix A). Somalwar et al. (2023b) found that the 2023 light curve was fast-evolving and red, along with the fact that the optical spectra obtained 36 days after the peak was consistent with Type Ia SN observed roughly 30 days post-peak. To the best of our knowledge, this is the first and only event where a TDE was followed by a SN in the same position. As discussed by Somalwar et al. (2023b), this scenario should be taken into account when identifying pTDEs and repeating TDEs.
The light curves for the two flares are plotted in Fig. 6(f). It is evident that the shape of the outbursts is very different. This is more evident in the bottom panel of Fig. 6(f), where the two flares are completely inconsistent with one another. The peak of the SN outburst is much sharper than the TDE one. Moreover, the rise and decay time scales are much faster in the former compared to the latter.
C.4 AT 2020vdq
This TDE was part of the Yao et al. (2023) sample. The first flare was detected in 2020, while the second flare was observed in 2023 (Charalampopoulos et al. 2023). Somalwar et al. (2023b) studied AT 2020vdq’s repeating flare, which also appeared in the sample of Somalwar et al. (2023a), where a radio counterpart detection was reported, 1.4 years after the 2020 flare peak. There is no available spectrum from the time of the first flare, however the spectrum taken during the 2023 flare is consistent with the TDE-He class. Somalwar et al. (2023b) conclude that only the pTDE scenario is likely, excluding the scenario that the two flares could be caused by two separate events.
However, they also state that the host galaxy of AT 2020vdq is a post-starburst galaxy (see their Table 8), meaning that the rate of TDEs occurring are boosted. Using the TDE rate in Yao et al. (2023), which is , and the fact that in post-starburst galaxies the TDE rate is enhanced by one or two factors of magnitude, Somalwar et al. (2023b) give a probability of 30% of two independent TDEs occurring in the same galaxy from the 33 TDEs in the Yao et al. (2023) sample. This leads to the conclusion that the two flares observed in AT 2020vdq could have been caused by two separate events.
Additionally, the two flares have different rise times and decay times. This disparity is clear in the bottom panel of Fig. 7(b). The second flare has a sharper rise and decay time, drawing similarities to the case of AT 2021mhg. Moreover, the second peak is more than two times higher than that of the first flare. There is also cooling during the second flare, while the first doesn’t show much evolution in the temperature (Somalwar et al. 2023b).
C.5 AT 2019azh
This TDE shows a lot of similarities with AT 2024pvu, in the sense that the initial flare was also observed by the CRTS survey, while the second was detected by another optical survey, in this case ZTF and Gaia surveys, roughly 14.5 years later. The peak for the first flare was not recorded, probably due to the seasonal gap. Nonetheless, the peak of the second flare was observed in March of 2019. Hinkle et al. (2021) have reported the existence of the first flare, but as they discuss, there is only photometric data available from one wave band, which cannot confirm or deny whether the outburst was caused by a TDE or some other transient. This is the same as for AT 2024pvu, where the initial flare was observed only in one waveband and no spectrum was taken. Hinkle et al. (2021) also placed the two flares on top of each other (see their Figure 3). Although there is limited data for the first flare, the two flares seem to decay in a similar manner, further showcasing the resemblance between repeating flares.
However, unlike AT 2024pvu, there are a lot of information about the host galaxy of AT 2019azh, KUG 0810+227. Hinkle et al. (2021) were able to rule out strong AGN activity using archival data of KUG 0810+227. Nevertheless, they do not exclude the presence of a weak or low-luminosity AGN (LLAGN; Tozzi et al. 2006; Marchesi et al. 2016; Liu et al. 2017; Ricci et al. 2017). Moreover, they classify the host as a post-starburst galaxy, which as we previously discussed for AT 2020vdq, has boosted TDE rates, meaning that the initial flare could have indeed been a separate TDE.
C.6 AT 2018fyk
This TDE was first studied in Wevers et al. (2019). It remained X-ray and UV bright for at least a period of 500 days post discovery. There was another detection in the X-rays, 1216 days after the initial discovery (Wevers et al. 2023), although not as bright as the first flare (factor lower than the first flare). Wevers et al. (2023) attributed the re-brightening to the partial disruption of a star that was a part of a binary star system. There is also a notable bump in the optical light curve obtained by the Gaia survey (see Fig. 11(a)). The first outburst was observed in 2018, while the re-brightening was detected in 2021. Similar behaviour can be observed in Fig. 11(b), where we plot the catalogue X-ray data (see Sect. B.5).
@@ Appendix D Repeating flare timescale comparison
To quantitatively investigate the similarity of the flares in the repeating flare sample, we estimate the rise times, decay times, peak flux and duration (metrics) for the non-confirmed TDE flares as well (i.e. first flare of AT2019teq, first flare of AT2022agi, first flare of AT2019azh, first flare of AT2024pvu, second flare of AT2021mhg). We note that we were unable to compute the rise time of AT2021mhg’s second flare and the decay time of the first flare of AT2019teq due to substantial gaps in the respective light curves. Additionally, we exclude AT2018fyk because the second flare is very faint in the optical band. For AT2022exr and AT2021uvz, which are the double peak flares, as well as AT2020acka, we cannot evaluate the shape of their flares since they are overlapping.
Figure 12 shows the metrics of the first flare vs the metrics of the second flare. Almost in all cases (except AT2019teq rise time and AT2024pvu in all timescales), the repeating flares are consistent within uncertainty. This was expected for AT2019teq, since the rise of the first flare is much faster than the second flare (TDE).

### v2

@@ B.3.3 Spectral extraction
The source spectra were extracted with the evselect task from circular regions centred at the X-ray source coordinates with radii equal , while background spectra were extracted from annuli centred at the X-ray source coordinates, with inner and outer radii equal to and , respectively, excluding nearby detected X-ray sources. The corresponding arf and rmf files were generated with the rmfgen and arfgen tasks to take into account time and position-dependent EPIC responses. Again, inner regions of high pileup were estimated and excluded using the epatplot task through the distortion of pattern distribution, following the procedure explained in the SAS Data Analysis Threads282828https://www.cosmos.esa.int/web/xmm-newton/sas-thread-epatplot.
@@ B.4 X-Ray spectral fitting
Spectral fitting was performed with the Sherpa292929http://cxc.harvard.edu/sherpa modelling and fitting application (Freeman et al. 2001) in the energy range for Swift-XRT and XMM-Newton-EPIC spectra, and in the energy range for Chandra-ACIS spectra, adopting Gehrels weighting (Gehrels 1986). Source spectra were binned to a minimum of 20 counts/bin to ensure the validity of statistics. In addition, for the EPIC spectra we excluded from the spectral fitting the band due to variable Al K lines, and fitted simultaneously MOS1, MOS2 and PN spectra, if available.
For the spectral fitting we used two different models: (1) a model comprising an absorption component fixed to the Galactic value and a power-law with slope and (2) a model comprising an absorption component fixed to the Galactic value and a black body with temperature . In order to select the best fit model we made use of the Akaike information criterion (AIC, see for example Liddle 2007):
where is the number of model parameters, so we selected as the best-fit model the one that provides the lower value of AIC.
B.5 X-ray catalogues
@@ B.5.1 XMM catalogue
We cross-matched the 4XMM-DR13 catalogue with our catalogue of TDEs using a search radius. Since the fluxes reported in 4XMM-DR13 are derived from the count rates assuming a model comprising a power-law with slope 1.7 and a Galactic absorption of (Webb et al. 2020), for uniformity we converted the count rates for each detector to observed intrinsic fluxes using PIMMS303030https://heasarc.gsfc.nasa.gov/docs/software/tools/pimms.html, assuming a model comprising an absorption component fixed to the Galactic value and a black body with a temperature of . The fluxes for each detector and their respective uncertainties were there converted in mean flux and uncertainty using the same procedure illustrated in Sect. B.3.2.
@@ B.5.2 EROSITA
We cross-matched the eRASS1 catalogue with our catalogue of TDEs using a search radius The fluxes reported in eRASS1 are derived from the count rates assuming a model comprising a power-law with slope 2.0 and a Galactic absorption of (Merloni et al. 2024), and therefore we converted the count rates reported in the catalogue to observed intrinsic fluxes using PIMMS, again assuming a model comprising an absorption component fixed to the Galactic value and a black body with a temperature of .
For the TDE sources for which we did not find an eRASS1 counterpart, we obtained eRASS1 upper limits313131https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/UpperLimitServer_dr1/. The upper limit eRASS fluxes were again converted in fluxes using PIMMS323232For the eRASS upper limits in the TDECat GitHub page, we provide the time range covered by the survey, since the exact exposure dates are not provided by the eRASS1 catalogue..
@@ Appendix C Archival repeating flares
C.1 AT 2018fyk
@@ Appendix D Repeating flare timescale comparison
To quantitatively investigate the similarity of the flares in the repeating flare sample, we estimate the rise times, decay times, peak flux and duration (metrics) for the non-confirmed TDE flares as well (i.e. first flare of AT 2022agi, first flare of AT 2019azh, first flare of AT 2024pvu, second flare of AT 2021mhg). We note that we were unable to compute the rise time of AT 2021mhg’s second flare due to substantial gaps in the light curves. Additionally, we exclude AT 2018fyk because the second flare is very faint in the optical band. For AT 2022exr and AT 2021uvz, which are the double peak flares, as well as AT 2020acka, we cannot evaluate the shape of their flares since they are overlapping.
Figure 11 shows the metrics of the first flare versus the metrics of the second flare. Almost in all cases (except AT 2024pvu in all timescales), the repeating flares are consistent within uncertainty.